OpenStudy (anonymous):

any help will receive a fanning Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following: •Does the graph open up or down? How do you know? (2 points) •Explain whether the graph has a maximum or minimum point. (2 points) •Find the vertex and x-intercepts of the graph. (3 points) next one is [9.01], [9.02] Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) How are the graphs alike and different.

OpenStudy (anonymous):

i get it then you can be dakohta fanning !

OpenStudy (anonymous):

lol be able to helkp?

OpenStudy (anonymous):

make up anything like say \(y=x^2+4x-3\) or anything you like

OpenStudy (anonymous):

in my example, since the leading coefficient is positive (it is one) it opens up

OpenStudy (anonymous):

since it opens up, it has a MINIMUM but no maximum 

OpenStudy (anonymous):

first coordinate of the vertex is always \(-\frac{b}{2a}\) which i my example is \(-\frac{4}{2\times 1}=-2\)

OpenStudy (anonymous):

second coordinate is what you get if you replace \(x\) by the first coordinate \[y=(-2)^2+4\times (-2)-3=4-8-3=-7\] so in my example it is \((-2,-7)\)

OpenStudy (anonymous):

soo confused

OpenStudy (anonymous):

oh and you need the \(x\) intercepts, so i picked a bad example as this one does not factor you can use the quadratic formula to find them but i find it easier to "complete the square" \[x^2+4x-3=0\\ x^2+4x=3\\ (x+2)^2=3+4=7\\ x+2=\pm\sqrt7\\x=-2\pm\sqrt7\]

OpenStudy (anonymous):

ok lets back up and go slow

OpenStudy (anonymous):

Create your own unique quadratic equation in the form y = ax2 + bx + c. Use complete sentences and show all work to determine the following:

OpenStudy (anonymous):

pick something anything that looks like \(y=ax^2+bx+c\) i.e. pick an \(a\) and a \(b\) and \(c\) make it easy by picking \(a=1\) and a nice even number for \(b\) let me know what you choose

OpenStudy (anonymous):

im so confuzzle

OpenStudy (anonymous):

did you pick some numbers? just pick 3 make one of them 1 and one of them even